integrate the following integral
∫ tan^2(2x – 3) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int tan^{2}(2x-3)dx
I = \int (sec^{2}(2x-3)-1)dx
I = \int sec^{2}(2x-3)dx-\int dx
I = tan(2x-3).\frac{1}{2}-\int dx
I = \frac{tan(2x-3)}{2}-x+c